OpenStudy (anonymous):

If paint cost $24 per gallon, how much would it cost per liter?

OpenStudy (anonymous):

\[\frac{24}{\text{gallon}}*\frac{\text{gallon}}{\text{liter}} \]

OpenStudy (anonymous):

\[\frac{24}{3.78541}=\text{$\$$6}.34013 \text{ per } \text{liter} \]